a train has a final philosophy of 110 km / H it accelerated for 36s
at 0.50 m / s what was its initial velocity
Answer #1

Sol 11oXS (8 SSO 191 ms. a= 0.50 m/ 2 0.50 m/2 ; t = 36 soc. - utat = u + 366 0.5 - le 550 - 18 = 12.56 me. Initial velocity